A Complete Bibliography of ACM Transactions on Computing for Healthcare (HEALTH)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

30 April 2024 Version 1.06

Title word cross-reference

2 [KGV21]. N [SCJ22].

19 [DL22, KDD^{+}22, San21].

2 [BCPM^{+}23, KX21].

5DoF [SMSV20].

Ability [EPM^{+}20]. ABIPA [WWF^{+}23]. Abnormality [TJH^{+}20]. Acceleration [CMS24]. Accelerometer [WWF^{+}23, ZAK^{+}23]. Accelerometer-Based [WWF^{+}23]. Accurate [ZAK^{+}23]. Across [CK20]. Actionable [DBJ23]. Active [CWT23, CM21]. Activities [AGJ21, DMWZ24]. Activity [ALD^{+}22, ABGO23, JBBF21, RFJKB21, RHK^{+}22, SCL^{+}23, TP24, VGS^{+}20, WWF^{+}23]. Adaptation [KBQvdS23]. Adapting [NKP^{+}22]. Adaptive [BCPM^{+}23, DCA23, WYZ^{+}20]. Adaptively [SCJ22]. Adolescent [WWF^{+}23]. Adoption [BFO^{+}21]. Adults [EPM^{+}20, VGS^{+}20]. Adversarial [ASA23, HDM^{+}20]. Advertising [YT20]. Affecting [FJK24]. Agent [CPP^{+}22]. Agnostic [YMG^{+}21]. aided [BYL^{+}23]. Alarm [PRB^{+}22]. Algorithm [AKMJ21]. All-textile [RHK^{+}22]. Alzheimer [BFO^{+}21]. Ambulatory [KMM^{+}22]. Analysis [ALD^{+}22, ABGO23, cJCCGW21, CPP^{+}22, KDD^{+}22, SHL^{+}22, YCW^{+}24].

Analytics [GLV22, HSK^{+}21, SOEM22]. Analyzing [ASL^{+}21, LYH^{+}23, TND^{+}24]. Anemia [GDUW22]. Anemia-Polycythemia [GDUW22]. Angiography [TVDB22]. Annotation [EWKM22]. Answering [BPDB21]. Anxiety [SOEM22]. Aortic [FLSC^{+}22]. Apnea [KNP^{+}21, RM22]. Application [ALD^{+}22, DMWZ24, GLV22]. Applications [FJK24, FHS22, JPS22, KMK^{+}21, KJBD21]. Applied [TJH^{+}20]. Approach [AAM^{+}24, BV20, CWT23, CM21, GDUW22, GPBB22, JJU^{+}22, NSA^{+}23, SBR^{+}23]. Approaches [RBS21]. Architectures [FLSC^{+}22]. Area [JJU^{+}22]. ARIMA [WWF^{+}23]. ARIMA-Based [WWF^{+}23]. Arrhythmia [KMM^{+}22]. Artifact [ZAK^{+}23]. Artificial [cJCCGW21, vBAS^{+}21]. Aspects [HSZ^{+}22]. Assessing [NP20]. Assessment [BCPM^{+}23, BYL^{+}23, GS22, KMK^{+}21, NSA^{+}23, RM22, RLC21, SOEM22, SCSEH22]. Assistance [JCG^{+}20]. Assistants [CM21]. Assisted [AAA^{+}20, KIT22, KPPA23]. Assisting [SQHS22]. Association [AKMJ21]. Attacks [NSRU21]. Attention [DL22, SDS^{+}21, TKK21]. Attention-Based [SDS^{+}21, DL22]. Attention-Gated [TKK21]. Augmentation [LHFC^{+}24, YLZ23]. Augmented [FHS22]. Augmenting [HDM^{+}20]. Authentication [CBF^{+}23]. Authoritative [HK23]. Automated [KZK^{+}21, SJMH20]. Automatic [BBMD^{+}22, BAL^{+}22, FLSC^{+}22, RBS21, SLT^{+}24, ZSL^{+}22]. Availability [KCTTM^{+}21]. Aware [AAA^{+}20, KMM^{+}22, XZC^{+}22, YMG^{+}21, SBR^{+}23].

Back [GLH^{+}23]. Barriers [BFO^{+}21]. Based [BKS20, CSSE22, FLSC^{+}22, KP23, San21, SDS^{+}21, WMKS21, WYM20, WYZ^{+}20, WHYM21, WWF^{+}23, AMN^{+}22, BPR^{+}22, DL22, FWWC23, NSA^{+}23, SLT^{+}24, SCSEH22, SAW22, TP24, YLZ23, ZAK^{+}23]. Bayesian [BV20, HK23, KDD^{+}22]. Beat [AAM^{+}24]. Behavior [ASA23, CSSE22, GLH^{+}23, SCSEH22, WOM^{+}21]. Behavioral [KDD^{+}22]. Behaviors [NP20]. Behaviour [KCTTM^{+}21]. Being [MFMF^{+}20, NP20, KCTTM^{+}21, BKS^{+}22]. Benefits [RFJKB21]. between [TVDB22]. Bias [MCM^{+}22, TJH^{+}20]. BioLumin [EWKM22]. Biomedical [BPDB21, DV22, EWKM22, GTC^{+}22]. Biometric [SAW22]. Bluetooth [RBS21]. Body [HRW^{+}20, JJU^{+}22]. Brain [CSSE22, RHK^{+}22, SMSV20]. Bridging [TVDB22].

Can [SPE^{+}21]. Cancer

[CMS24, LHFC^{+}24, YT20]. Capabilities [San21]. Cardiovascular [HSK^{+}21]. Care [GLV22]. CarePre [JCG^{+}20]. Case [KCTTM^{+}21, KDD^{+}22, TVDB22, vBAS^{+}21]. Case-study [TVDB22]. Catheter [SMSV20]. Causal [HK23, KBQvdS23]. CAVE [EPM^{+}20]. centered [CM21]. Challenges [HSK^{+}21, JPS22, MKE^{+}20]. Change [CSSE22, KDD^{+}22]. Changes [WKS^{+}21]. Characteristics [MKE^{+}20]. Chatbots [CM21, SPE^{+}21]. Checking [CK20]. Chronic [GLH^{+}23, WOM^{+}21]. Class [YMG^{+}21]. Class-Agnostic [YMG^{+}21]. Classification [CM21, MCM^{+}22, TJH^{+}20, WKT24]. Classifier [BAL^{+}22, NKP^{+}22, NKP^{+}22]. Classifiers [PRB^{+}22]. Clean [FBS^{+}22]. Clinical [BBMD^{+}22, CM21, HDD^{+}23, JCG^{+}20, SHL^{+}22, SCSEH22]. Cluster [KZK^{+}21]. Cluster-Driven [KZK^{+}21]. Clusters [KZK^{+}21]. Co [XZC^{+}22]. Co-Located [XZC^{+}22]. Cognitive [BAL^{+}22]. Cohort [KIT22]. Collected [ASL^{+}21, LVW^{+}20]. Collection [KMM^{+}22]. College [BKS^{+}22]. Colonoscopy [vBAS^{+}21]. Combinations [NP20]. Combining [WKT24]. Commodity [MPL^{+}20, WYM20]. Community [KDD^{+}22]. Community-scale [KDD^{+}22]. Comparing [EPM^{+}20]. Complex [AGJ21]. Comprehensive [FJK24, HSZ^{+}22, SO22]. Compressed [SAW22]. Compressed-Domain [SAW22]. Compressive [FWWC23]. Computational [DV22, RLC21, SO22]. Computer [KIT22]. Computer-Assisted [KIT22]. confidence [PRB^{+}22]. Conformance [CK20]. Conjugate [CMS24]. Consent [VYJ21]. Consultations [APH^{+}20]. Consumer [ALF^{+}23]. Consumer-Grade [ALF^{+}23]. Contact [RBS21, WHYM21]. Contact-Free [WHYM21]. Context [AGJ21, AAA^{+}20, KMM^{+}22, LKHU22, BKS^{+}22, XZC^{+}22]. Context-Aware [AAA^{+}20, KMM^{+}22, XZC^{+}22]. Context-Enhanced [LKHU22]. Continuous [MPL^{+}20, vBAS^{+}21]. Contrastive [LHFC^{+}24]. Conversational [CPP^{+}22]. Convolutional [BPR^{+}22, KGV21, ZSL^{+}22]. Convolutions [TKK21]. Coronary [TVDB22]. Correlation [CWR^{+}21]. Cost [CWR^{+}21, CWT23]. Cost-Effective [CWR^{+}21, CWT23]. Counterfactuals [DBJ23]. Counts [SCL^{+}23]. COVID [DL22, KDD^{+}22, San21]. COVID-19 [DL22, KDD^{+}22, San21]. Creating [CM21]. Crisis [SO22]. Critical [SOEM22]. Cross [TJH^{+}20]. Cross-Dataset [TJH^{+}20]. CSI [WYM20]. CSI-Based [WYM20]. CT [TVDB22]. Customized [HRW^{+}20]. CycleGAN [ZAK^{+}23].

D [KGV21]. Daily [AGJ21, DMWZ24]. Data [AGJ21, BV20, CPP^{+}22, DCA23, GKKS20, GPBB22, HCZ21, KMK^{+}21, KCTTM^{+}21, LVW^{+}20, LHFC^{+}24, MFMF^{+}20, MKE^{+}20, MCM^{+}22, NKP^{+}22, PRB^{+}22, SCL^{+}23, TND^{+}24, WKS^{+}21, YLZ23, YCW^{+}24]. Data-driven [AGJ21, DCA23]. Dataset [TJH^{+}20]. Datasets [MNMC21]. De-identification [LKHU22]. Decision [BFO^{+}21, JCG^{+}20, MJPMC22]. Decision-Support [BFO^{+}21]. Deep [CWT23, CM21, DCA23, GPBB22, KPPA23, MNMC21, NSA^{+}23, SDS^{+}21, WOM^{+}21, WKT24, ZSL^{+}22]. Defenses [NSRU21]. Delivery [San21]. Dementia [BAL^{+}22, KAG^{+}22]. Depression [SLT^{+}24]. Depressive [RFJKB21]. Depth [XZC^{+}22]. Designing [KCTTM^{+}21, vBAS^{+}21]. Detecting [CSSE22]. Detection [AGJ21, GSG^{+}22, GDUW22, KNP^{+}21, LBHS22, MPL^{+}20, SJMH20, WOM^{+}21, WKS^{+}21]. Deterioration [LVW^{+}20]. Development [Thi22]. Devices [MPK20, WYZ^{+}20]. Diabetes [BCPM^{+}23]. DiaFocus [BCPM^{+}23]. Diagnosis [cJCCGW21, LHFC^{+}24]. Diagnostics [MPK20]. Dicrotic [SBR^{+}23]. Dictionary [SAW22]. Different [HSZ^{+}22]. Differentially [AAF^{+}22, AKMJ21]. Digital [Thi22]. Dim [WMKS21]. Discharge [CK20]. Disease [ASL^{+}21, BFO^{+}21]. Disorders [HSK^{+}21]. Displays [HRW^{+}20]. Dissection [FLSC^{+}22]. Distance [APH^{+}20]. Distortions [GSG^{+}22]. distribution [GSG^{+}22]. Diversifying [BMC^{+}21]. DL [HDM^{+}20]. Domain [CSSE22, GTC^{+}22, JPS22, KBQvdS23, SAW22]. Domain-Specific [GTC^{+}22]. Driven [KZK^{+}21, SAW22, AGJ21, DCA23, HDD^{+}23]. Drug [TKK21]. During [RHK^{+}22]. Dynamic [AAA^{+}20, SO22].

Early [BAL^{+}22, KAG^{+}22, WKS^{+}21].

Eating [BKS^{+}22]. ECG

[cJCCGW21, KMM^{+}22, SAW22]. ECG-based [SAW22]. Edge [AAA^{+}20, RM22]. Edge-Assisted [AAA^{+}20]. Editorial [SL20]. EEG [GP24]. Effective [CWR^{+}21, CWT23, LHFC^{+}24].

Effects [KBQvdS23]. Efficient [JJU^{+}22, KMK^{+}21, LCLS20, BPR^{+}22, DCA23]. EHR [GPBB22]. Elderly [WKS^{+}21]. Electronic [LYH^{+}23, VYJ21, YCR^{+}24]. Elements [GPBB22]. Eligibility [CM21]. Embedding [BPR^{+}22]. Emotion [GSG^{+}22, WYZ^{+}20]. Emotions [GSG^{+}22]. EmotionSense [WYZ^{+}20]. Empirical [CWR^{+}21]. Empowering [CPP^{+}22]. Enabling [CWR^{+}21]. Encoding [WKT24]. End [NKP^{+}22]. End-User [NKP^{+}22]. Energy [BPR^{+}22, DCA23, JJU^{+}22, RBS21]. Energy-efficient [BPR^{+}22, DCA23]. Enhanced [LKHU22, MFMF^{+}20]. Enhancing [LHFC^{+}24]. Ensemble [BMC^{+}21, MNMC21, TP24]. Entities [BBMD^{+}22]. Entity [BPDB21]. Environmental [GSG^{+}22]. Epilepsy [HDM^{+}20]. Error [AAM^{+}24]. Estimation [AAM^{+}24, HDD^{+}23, LBHS22, WMKS21]. Evaluating [CM21, PRB^{+}22]. Evaluation [FLSC^{+}22, TJH^{+}20]. EVD [SMSV20]. Everyday [GP24]. Evolution [BMC^{+}21]. Examining [BKS^{+}22]. Exergaming [EPM^{+}20]. Experience [EWKM22]. Expertise [HK23]. Experts [HDD^{+}23, SPE^{+}21]. Explaining [SHL^{+}22]. Explanations [BMC^{+}21]. Exploiting [CWR^{+}21]. Exponential [CMS24]. Expression [PR22]. Extensive [TJH^{+}20]. Extracting [TKK21]. Extraction [BBMD^{+}22, CPP^{+}22, DL22]. Extreme [TP24]. Eye [RHK^{+}22].

Facial [PR22, SLT^{+}24]. Factors [FJK24, HSK^{+}21]. Fall [LBHS22]. Fast [cJCCGW21]. Feasibility [LVW^{+}20]. Feature [GS22, GKKS20]. Features [WMKS21]. Federated [JPS22]. Film [HRW^{+}20]. Flexibility [AMN^{+}22]. Flexible [BV20, HRW^{+}20]. Focused [GKKS20]. Food [MJPMC22, BKS^{+}22, YMG^{+}21]. form [TVDB22]. Framework [HCZ21, JBBF21, YSV23]. Frameworks [KJBD21]. Free [TVDB22, WHYM21]. Free-form [TVDB22]. Function [ALF^{+}23]. Fundus [GS22, KP23]. Fusion [LHFC^{+}24, SCSEH22, XZC^{+}22]. Future [KP23, PR22].

Gait [SHL^{+}22]. Game [FHS22]. GAN [FWWC23, YLZ23]. GAN-based [FWWC23]. Gap [TVDB22]. Gastrointestinal [TJH^{+}20]. Gated [TKK21]. GeCoAgent [CPP^{+}22]. Genderization [MCM^{+}22]. Generalized [KMK^{+}21]. Generating [HCZ21]. Generation [ASA23, AAF^{+}22]. Generative [ASA23, AAF^{+}22]. Genome [AKMJ21]. Genome-wide [AKMJ21]. Genomic [CPP^{+}22]. Gland [ZSL^{+}22]. Glaucoma [GS22]. GPU [CMS24]. Grade [ALF^{+}23]. Graph [DL22, TKK21, YCR^{+}24]. Growth [MPMP21]. Guaranteed [CWT23]. Guidance [BPDB21, MPK20]. Guidelines [CK20]. Gymnastics [KMK^{+}21].

Hadamard [TP24]. Hand [FBS^{+}22]. Hardware [MPMP21, MPL^{+}20]. Head [ASL^{+}21]. Headset [GP24]. Health [AMN^{+}22, BCPM^{+}23, CWR^{+}21, CWT23, CSSE22, FJK24, FWWC23, GLV22, HCZ21, JJU^{+}22, KX20, KX21, MFMF^{+}20, MKE^{+}20, NKP^{+}22, NP20, SPE^{+}21, Thi22, WKS^{+}21, YLZ23, YCR^{+}24]. Healthcare [CBF^{+}23, DBJ23, FHS22, GKKS20, JPS22, JBBF21, KP23, MPK20, NSRU21, SPE^{+}21, TND^{+}24]. Healthy [MJPMC22]. Heart [BPR^{+}22, NSA^{+}23]. Help [SPE^{+}21]. Heterogeneity [CMS24]. High [PRB^{+}22, SMSV20]. High-confidence [PRB^{+}22]. High-Precision [SMSV20]. HMD [EPM^{+}20]. Home [BKS20, KNP^{+}21, SCSEH22, WKS^{+}21, XZC^{+}22]. Home-Based [BKS20]. Hospital [APH^{+}20]. Human [ABGO23, MFMF^{+}20, SQHS22, TP24]. Hybrid [GDUW22]. Hypothyroidism [GLV22].

Identification [KIT22, SBR^{+}23, SAW22, LKHU22]. Image [YMG^{+}21, WKT24]. Images [GS22, KGV21]. Imaging [KP23]. Imaging-Based [KP23]. Immersive [EPM^{+}20, EWKM22]. Impact [SJMH20]. Impairment [BAL^{+}22]. Implications [Thi22]. Improving [HK23, KAG^{+}22, MFMF^{+}20, NP20]. Imputation [FWWC23]. IMU [KMK^{+}21]. In-Home [WKS^{+}21, XZC^{+}22]. iNAP [GDUW22]. Inaugural [SL20]. Indicators [AMN^{+}22]. Indoor [GSG^{+}22]. Infection [BKS20]. Inferring [SCL^{+}23]. Influences [MJPMC22]. Infomax [YCR^{+}24]. Information [AMN^{+}22, SQHS22, TKK21]. Innovations [MJPMC22]. Insights [CBF^{+}23]. Insoles [DMWZ24]. Inspired [BKS20]. Integration [WWF^{+}23]. Intelligence [vBAS^{+}21]. Intelligent [JCG^{+}20]. Inter [AAM^{+}24]. Inter-Beat [AAM^{+}24]. Interaction [MFMF^{+}20, TKK21]. Interactive [EWKM22]. Internet [BGL22, YT20]. Internet-of-Medical-Things [BGL22]. Interpretable [GPBB22, MCM^{+}22, YCW^{+}24]. Interpretation [TJH^{+}20]. Interval [AAM^{+}24]. Intervention [CSSE22]. Interventions [KCTTM^{+}21, KDD^{+}22]. Introduction [BGL22, DV22, KX20, KX21]. invasive [HSZ^{+}22]. IoMT [GDUW22]. IoT [BKS20, JBBF21]. IoT-Inspired [BKS20]. Issue [BGL22, DV22, KX20, KX21, SL20]. Issues [NSRU21].

Kernel [YCR^{+}24]. Keyphrase [DL22]. Knowledge [BPDB21, KBQvdS23].

Labels [TKK21]. Language [GTC^{+}22].

Learned [YSV23]. Learner [LHFC^{+}24]. Learning [ABGO23, BMC^{+}21, CWT23, CM21, DCA23, GPBB22, JPS22, KAG^{+}22, KNP^{+}21, KPPA23, MNMC21, NSA^{+}23, SQHS22, SHL^{+}22, SAW22, TP24, TJH^{+}20, WOM^{+}21, WKT24, YCR^{+}24, YT20]. Learning-assisted [KPPA23]. Learning-based [NSA^{+}23]. length [ALD^{+}22]. Lengths [LCLS20]. Lessons [YSV23]. Leveraging [AGJ21, KDD^{+}22]. Life [MKE^{+}20]. Light [WMKS21]. Lightweight [LHFC^{+}24]. Limited [San21]. Literature [DL22, FHS22, VYJ21, YSV23]. Liver [LHFC^{+}24]. Living [AGJ21, DMWZ24]. Located [XZC^{+}22]. Long [BCPM^{+}23, MKE^{+}20]. Long-Term [BCPM^{+}23]. Longitudinal [AMN^{+}22, BV20, CBF^{+}23, YCW^{+}24]. Loss [SLT^{+}24]. Low [GLH^{+}23, RBS21]. Lung [ALF^{+}23].

Machine [BYL^{+}23, KAG^{+}22, KNP^{+}21, SQHS22, SHL^{+}22, TP24, TJH^{+}20]. Machine-aided [BYL^{+}23]. Maintaining [MCM^{+}22]. Mammographic [KGV21]. Management [BCPM^{+}23, RFJKB21]. Mapping [GS22]. Markers [SCSEH22, vBAS^{+}21]. MARS [SQHS22]. Mask [RHK^{+}22]. Masks [ALF^{+}23]. Matching [HDD^{+}23, SCJ22]. mCardia [KMM^{+}22]. Meals [SJMH20]. Measurements [GKKS20]. Measuring [ALF^{+}23]. Media [GLV22, RLC21]. Medical [AAF^{+}22, BV20, BGL22, DL22, HRW^{+}20, HK23, KMK^{+}21, LYH^{+}23]. Medical-Order [LYH^{+}23]. Melatonin [WMKS21]. Mental [BKS^{+}22, RLC21, SPE^{+}21, WKT24]. Method [FWWC23, RM22]. Methodology [CK20]. Methods [DV22, KAG^{+}22, LYH^{+}23, RLC21]. Metrics [TJH^{+}20]. mHealth [KJBD21, YSV23]. mHealth-Literature [YSV23]. Microscopic [EWKM22]. Mild [BAL^{+}22]. Minimalist [RM22]. Mining [LYH^{+}23, LCLS20]. Mitigation [MCM^{+}22]. Mixed [EPM^{+}20, EWKM22, FHS22]. Mixture [HDD^{+}23]. Mobile [FJK24, KCTTM^{+}21, KJBD21, KDD^{+}22]. Modal [WKT24]. Modality [LHFC^{+}24]. mode [BYL^{+}23]. Model [AAM^{+}24, BMC^{+}21, CMS24, GTC^{+}22, MFMF^{+}20, San21, SDS^{+}21]. Modeling [PR22]. Modelling [BV20, KZK^{+}21]. Models [KCTTM^{+}21, KBQvdS23, SO22, SHL^{+}22]. Modern [NSRU21]. Monitor [KDD^{+}22]. Monitoring [BYL^{+}23, BPR^{+}22, CWR^{+}21, FBS^{+}22, HSZ^{+}22, JJU^{+}22, JBBF21, KNP^{+}21, LVW^{+}20, RM22, WYM20, WHYM21, XZC^{+}22]. Mood [KZK^{+}21, BKS^{+}22]. Motion [SJMH20, ZAK^{+}23]. Movement [GLH^{+}23]. mSIMPAD [LCLS20]. Multi [BYL^{+}23, CSSE22, KGV21, LHFC^{+}24, WKT24, WKS^{+}21]. Multi-Domain [CSSE22]. Multi-Modal [WKT24]. Multi-Modality [LHFC^{+}24]. Multi-mode [BYL^{+}23]. Multi-Sensor [WKS^{+}21]. Multi-task [KGV21]. Multi-view [KGV21]. Multimodal [AMN^{+}22, LVW^{+}20, SOEM22, SCSEH22]. Multimodal-multisensor [SOEM22]. Multiple [LCLS20]. multisensor [SOEM22]. MVCT [ZSL^{+}22]. My [NKP^{+}22, NKP^{+}22].

Named [BPDB21]. Natural [GTC^{+}22]. Navigation [HRW^{+}20]. Nearables [AGJ21]. Needle [HRW^{+}20]. Nested [BBMD^{+}22]. Net [FLSC^{+}22]. Network [cJCCGW21]. Networks [ASA23, AAF^{+}22, ABGO23, BPR^{+}22, HK23, JJU^{+}22, KGV21, San21, SCL^{+}23, YCW^{+}24, ZSL^{+}22]. Neural [AAF^{+}22, ABGO23, cJCCGW21, KGV21, SCL^{+}23, YCW^{+}24, ZSL^{+}22]. NHS [Thi22]. NLP [DV22]. Noisy [SLT^{+}24]. Non [cJCCGW21, GSG^{+}22, HSZ^{+}22, ZAK^{+}23].

Non-accelerometer-based [ZAK^{+}23]. Non-invasive [HSZ^{+}22]. Non-targeted [GSG^{+}22]. Non-Uniform [cJCCGW21]. NonInvasive [GDUW22]. Nonparametric [BV20]. Notch [SBR^{+}23]. Notes [MCM^{+}22]. Novel [AAM^{+}24, KZK^{+}21, MPMP21]. Number [Thi22].

Obesity [GPBB22]. Obstructive [RM22]. OCT [GS22]. Office [KCTTM^{+}21]. Older [VGS^{+}20]. On-Body [HRW^{+}20]. Online [AKMJ21, RLC21]. Onset [WMKS21]. Ontologies [HK23]. Open [Thi22]. Opioid [SO22]. Opportunities [BFO^{+}21, HSK^{+}21]. Optimal [San21]. Order [LYH^{+}23]. Organ [SCJ22]. Organizations [CBF^{+}23]. Out-of-distribution [GSG^{+}22]. Outcomes [HDD^{+}23]. Outcomes-driven [HDD^{+}23]. Outpatients [LVW^{+}20].

Pain [BMC^{+}21, GLH^{+}23, WOM^{+}21].

Paradigm [CBF^{+}23]. Parkinson [ASL^{+}21]. Parotid [ZSL^{+}22]. Part [KX21]. Passive [XZC^{+}22]. Passively [AGJ21]. Past [PR22, YSV23]. Patient [HDD^{+}23, MCM^{+}22, PR22]. Patients [GLH^{+}23]. Pattern [LYH^{+}23]. Patterns [LCLS20, SCL^{+}23]. People [ASL^{+}21]. Perception [KZK^{+}21, MFMF^{+}20]. Perceptions [SPE^{+}21]. Performance [MCM^{+}22]. Person [SPE^{+}21]. Personal [BCPM^{+}23, HCZ21, JBBF21]. Personalized [CBF^{+}23, GLV22]. Perspective [TND^{+}24]. Pervasive [LBHS22]. Phenotyping [HDD^{+}23]. Phrase [DL22]. PhyMask [RHK^{+}22]. Physical [ALD^{+}22, VGS^{+}20, WWF^{+}23]. Physics [SBR^{+}23]. Physics-aware [SBR^{+}23]. Physiological [BYL^{+}23, MPL^{+}20, RHK^{+}22]. Physiotherapy [APH^{+}20]. Physiowise [SBR^{+}23]. Picture [FJK24]. Pipeline [JPS22]. Placement [SMSV20]. Point [KDD^{+}22]. Polycythemia [GDUW22]. Population [CWR^{+}21, CWT23, FWWC23]. Pose [ASL^{+}21, LBHS22]. PPG [BYL^{+}23, BPR^{+}22, NSA^{+}23, ZAK^{+}23]. PPG-based [BPR^{+}22]. Practice [KIT22]. Precision [SMSV20]. Predicting [BMC^{+}21]. Prediction [BKS20, GPBB22, HDM^{+}20, KCTTM^{+}21, MPMP21, SCSEH22, SDS^{+}21, WWF^{+}23]. Prematurity [KPPA23]. Present [KP23, PR22]. preserving [JBBF21]. Pretraining [GTC^{+}22]. Privacy [JBBF21, NSRU21, TND^{+}24, VYJ21]. Privacy-preserving [JBBF21]. Private [AAF^{+}22, AKMJ21]. Processing [GTC^{+}22, SQHS22]. Professionals [SPE^{+}21]. Profiling [CWT23]. Prognosis [KAG^{+}22]. Programmability [MFMF^{+}20]. Programming [AAA^{+}20, PRB^{+}22]. Progression [RM22]. Protective [WOM^{+}21]. Put [HRW^{+}20].

Quality [BYL^{+}23, CWT23, NSA^{+}23]. Quality-Guaranteed [CWT23]. Quantifying [GLH^{+}23]. Question [BPDB21].

Radiology [TVDB22]. Randomized [TP24]. Rate [BPR^{+}22, NSA^{+}23, WHYM21]. Reality [EPM^{+}20, EWKM22, FHS22, GLH^{+}23]. Recognition [ABGO23, BPDB21, DMWZ24, GSG^{+}22, JBBF21, KZK^{+}21, SLT^{+}24, TP24, WYZ^{+}20]. Recognizing [AGJ21]. Recommendation [SCJ22]. Recommender [RFJKB21]. Recommending [NP20]. Reconfigurable [MPMP21]. Record [LYH^{+}23]. Records [YCR^{+}24]. Recurrent [SCL^{+}23, SDS^{+}21]. Reduced [AAM^{+}24]. Reducing [MCM^{+}22]. Referrals [BBMD^{+}22]. Regulation [Thi22]. Reinforcement [DCA23]. Relaxation [SLT^{+}24]. Relief [BMC^{+}21]. Remotely [ASL^{+}21]. Removal [ZAK^{+}23]. Reporting [TVDB22]. Representation [YCR^{+}24]. Representational [ABGO23]. Requirements [VYJ21]. Research [EWKM22]. Respiration [WHYM21]. Response [CSSE22]. Resting [SJMH20]. Retinopathy [KPPA23]. Retrieval [DL22]. Review [DMWZ24, FHS22, HSZ^{+}22, KJBD21, SOEM22, VGS^{+}20, VYJ21, YSV23]. Risk [HSK^{+}21, HDD^{+}23]. Robust [GSG^{+}22, HDM^{+}20, KCTTM^{+}21, LCLS20, LHFC^{+}24, RHK^{+}22]. ROP [KPPA23]. Routineness [AMN^{+}22].

Saliency [YMG^{+}21]. Saliency-Aware [YMG^{+}21]. Sampling [DCA23]. scale [KDD^{+}22]. Scales [WMKS21]. SCouT [DBJ23]. Screening [KMM^{+}22, KPPA23, YT20]. Security [CBF^{+}23, NSRU21, TND^{+}24, VYJ21]. Segmentation [FLSC^{+}22, YMG^{+}21, ZSL^{+}22]. Seizures [HDM^{+}20]. Selecting [KBQvdS23]. Selection [GKKS20]. Self [MKE^{+}20, RFJKB21, YCR^{+}24, YSV23, YSV23]. Self-management [RFJKB21]. Self-Supervised [YCR^{+}24]. Self-Tracking [MKE^{+}20, YSV23]. Sensed [AGJ21]. Sensing [AAA^{+}20, FBS^{+}22, HSK^{+}21, KCTTM^{+}21, KJBD21, KDD^{+}22, RHK^{+}22]. Sensor [AMN^{+}22, NKP^{+}22, WMKS21, WKS^{+}21, XZC^{+}22, YLZ23]. Sensor-Based [WMKS21, AMN^{+}22, YLZ23]. Sequence [ASA23, LYH^{+}23]. Sequential [LYH^{+}23]. Series [ALD^{+}22, LCLS20, YLZ23]. Settings [APH^{+}20]. Severity [RM22]. Shooting [EPM^{+}20]. Sign [WYM20]. Signal [BYL^{+}23, WKT24]. Signal-image [WKT24]. Signals [RHK^{+}22]. Signs [XZC^{+}22]. Similar [LCLS20]. Simulations [SO22]. Simulator [PR22]. Situ [AMN^{+}22]. Skill [KMK^{+}21]. Sleep [HSZ^{+}22, KNP^{+}21, RM22, RHK^{+}22]. Small [MNMC21]. Smart [BKS20, DMWZ24, KX20, KX21, SCSEH22, WYZ^{+}20]. Smartphone [WHYM21]. Smartwatch [CSSE22]. Smartwatch-Based [CSSE22]. Social [GLV22, RLC21, San21]. Software [Thi22]. SoK [TND^{+}24]. Sonar [WHYM21]. Sonar-Based [WHYM21]. Source [Thi22]. Space [GS22]. Spanish [BBMD^{+}22]. Sparse [KCTTM^{+}21, SCL^{+}23]. Spatiotemporal [CWR^{+}21, DBJ23]. Special [BGL22, DV22, KX20, KX21]. Specific [GTC^{+}22]. Spectral [cJCCGW21]. Speech [BAL^{+}22]. SpiroMask [ALF^{+}23]. SQA [BYL^{+}23]. Standard [CK20]. Stars [EPM^{+}20]. State [RLC21]. Status [WWF^{+}23]. Step [SCL^{+}23]. Strategies [San21]. Strategy [SLT^{+}24]. Streams [WKS^{+}21]. Stress [MPL^{+}20]. Structural [FBS^{+}22]. Structured [TVDB22]. Student [BKS^{+}22]. Studies [AKMJ21, KJBD21]. Study [AMN^{+}22, CWR^{+}21, CBF^{+}23, KCTTM^{+}21, KDD^{+}22, LVW^{+}20, TJH^{+}20, vBAS^{+}21, TVDB22]. Subsampled [TP24]. Successive [LCLS20]. Summaries [CK20, HCZ21]. Summarization [ALD^{+}22]. SummerTime [ALD^{+}22]. Supervised [YCR^{+}24]. Support [BFO^{+}21, SPE^{+}21, vBAS^{+}21]. Supporting [GLV22]. Survey [HSK^{+}21, MJPMC22, NSRU21, RBS21, RLC21, SO22]. Survival [SDS^{+}21]. Sustainable [FWWC23]. Symptoms [RFJKB21]. Synthesis [PR22]. Synthetic [ASA23, DBJ23]. System [BKS20, JCG^{+}20, KZK^{+}21, KMM^{+}22, LKHU22, WYZ^{+}20, YT20]. Systematic [DMWZ24, FHS22, KJBD21, VGS^{+}20, VYJ21]. Systems [AAA^{+}20, LYH^{+}23, MPMP21, NSRU21, PR22, VGS^{+}20].

Target [GKKS20]. Target-Focused [GKKS20]. targeted [GSG^{+}22]. Task [SAW22, KGV21]. Task-Driven [SAW22]. Tasks [SQHS22]. Technique [ZAK^{+}23]. Techniques [HSZ^{+}22]. Technological [MJPMC22]. Technologies [KX20, KX21]. Technology [APH^{+}20, BCPM^{+}23, RFJKB21, YSV23]. Temporal [BPR^{+}22, HCZ21]. Term [BCPM^{+}23, MKE^{+}20]. Text [SCSEH22]. Text-based [SCSEH22]. textile [RHK^{+}22]. Texts [AAF^{+}22]. Textual [MCM^{+}22]. There [HRW^{+}20]. Thin [HRW^{+}20]. Thin-Film [HRW^{+}20]. Things [BGL22]. Three [CBF^{+}23]. Time [ALD^{+}22, LCLS20, WMKS21, YLZ23]. Time-series [YLZ23]. Tiramisu [AAM^{+}24]. Tissue [CMS24]. Toilet [BKS20]. Tool [BFO^{+}21]. Top [SCJ22]. Top- [SCJ22]. Tracing [RBS21]. Tracking [MKE^{+}20, SMSV20, VGS^{+}20, YSV23]. Tract [TJH^{+}20]. Trained [NKP^{+}22]. Training [HDM^{+}20, KMK^{+}21]. Transfer [ABGO23]. Transformation [MPK20, TP24]. Transformation-based [TP24]. Transformers [DBJ23]. Treatment [CK20, KBQvdS23, MPK20]. Trend [YCW^{+}24]. Triage [KGV21]. Trials [CM21]. Truly [GP24]. TS [YLZ23]. TS-GAN [YLZ23]. Tumor [MPMP21]. Two [WMKS21]. Type [BCPM^{+}23].

U [FLSC^{+}22]. U-Net [FLSC^{+}22]. Unattended [KNP^{+}21]. Uncertainty [GKKS20]. Uniform [cJCCGW21]. Unlabeled [NKP^{+}22]. Unsupervised [DL22]. Urine [BKS20]. Usability [CBF^{+}23]. Use [APH^{+}20, FJK24, GP24]. User [CM21, CBF^{+}23, FJK24, NKP^{+}22, SAW22, TND^{+}24]. User-centered [CM21]. Using [ALF^{+}23, ASA23, AAF^{+}22, ABGO23, cJCCGW21, DCA23, FBS^{+}22, GSG^{+}22, GKKS20, HK23, KAG^{+}22, KZK^{+}21, KCTTM^{+}21, KGV21, KBQvdS23, LVW^{+}20, MNMC21, RBS21, RFJKB21, SQHS22, SMSV20, SAW22, WMKS21, WYM20, WKS^{+}21, XZC^{+}22, YT20, ZSL^{+}22, ZAK^{+}23]. UWB [XZC^{+}22]. UWB-Depth [XZC^{+}22].

Vaccination [San21]. Vaccine [San21]. Variability [NSA^{+}23]. Variable [ALD^{+}22]. Variable-length [ALD^{+}22]. Variants [LYH^{+}23]. Versus [AMN^{+}22]. Very [MKE^{+}20]. via [AAA^{+}20, BPDB21, DBJ23]. Vibes [FBS^{+}22]. Vibration [FBS^{+}22]. Video [APH^{+}20, SLT^{+}24]. Video-based [SLT^{+}24]. Videos [ASL^{+}21]. view [KGV21]. Views [SPE^{+}21]. Virtual [EPM^{+}20, FHS22, GLH^{+}23]. Visual [vBAS^{+}21]. Visualization [EWKM22, SMSV20]. Vital [WYM20, XZC^{+}22].

Walking [SJMH20]. WalkingWizard [GP24]. Washing [FBS^{+}22]. Wearable [APH^{+}20, AAA^{+}20, GP24, KX20, KX21, KJBD21, MPK20, VGS^{+}20, WYZ^{+}20]. Wearables [LVW^{+}20, MNMC21]. Weight [WWF^{+}23]. Weighted [SCJ22]. Well [KCTTM^{+}21, MFMF^{+}20, BKS^{+}22, NP20]. Well-Being [MFMF^{+}20, NP20, KCTTM^{+}21, BKS^{+}22]. Wellbeing [WKT24]. Wellness [MKE^{+}20]. While [MCM^{+}22]. wide [AKMJ21]. WiFi [WYM20]. Willingness [FJK24]. Wireless [JJU^{+}22]. Worker [KCTTM^{+}21]. Workers [AMN^{+}22]. Wrist [SJMH20].

Year [FWWC23]. Year-By-Year [FWWC23]. Years [YSV23].

References

Amiri:2020:CAS

[AAA^{+}20] Delaram Amiri, Arman Anzanpour, Iman Azimi, Marco Levorato, Pasi Liljeberg, Nikil Dutt, and Amir M. Rahmani. Context-aware sensing via dynamic programming for edge-assisted wearable systems. ACM Transactions on Computing for Healthcare (HEALTH), 1(2):7:1–7:25, March 2020. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3351286.

AlAziz:2022:DPM

[AAF^{+}22] Md Momin Al Aziz, Tanbir Ahmed, Tasnia Faequa, Xiaoqian Jiang, Yiyu Yao, and Noman Mohammed. Differentially private medical texts generation using generative neural networks. ACM Transactions on Computing for Healthcare (HEALTH), 3(1): 5:1–5:27, January 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3469035.

Arefeen:2024:IBI

[AAM^{+}24] Asiful Arefeen, Ali Akbari, Seyed Iman Mirzadeh, Roozbeh Jafari, Behrooz A. Shirazi, and Hassan Ghasemzadeh. Inter-beat interval estimation with tiramisu model: a novel approach with reduced error. ACM Transactions on Computing for Healthcare (HEALTH), 5(1):1:1–1:??, January 2024. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3616020.

An:2023:TLH

[ABGO23] Sizhe An, Ganapati Bhat, Suat Gumussoy, and Umit Ogras.

Transfer learning for human activity recognition using representational analysis of neural networks. ACM Transactions on Computing for Healthcare (HEALTH), 4(1):5:1–5:??, January 2023. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3563948.

Akbari:2021:DDC

[AGJ21] Ali Akbari, Reese Grimsley, and Roozbeh Jafari. Datadriven context detection leveraging passively sensed nearables for recognizing complex activities of daily living. ACM Transactions on Computing for Healthcare (HEALTH), 2(2): 12:1–12:22, April 2021. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3428664.

Aziz:2021:OAD

[AKMJ21] Md Momin Al Aziz, Shahin Kamali, Noman Mohammed, and Xiaoqian Jiang. Online algorithm for differentially private genome-wide association studies. ACM Transactions on Computing for Healthcare (HEALTH), 2(2):13:1–13:27, April 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3431504.

Amaral:2022:SVL

[ALD^{+}22] Kevin Amaral, Zihan Li, Wei Ding, Scott Crouter, and Ping Chen. SummerTime: Variablelength time series summarization with application to physical activity analysis. ACM Transactions on Computing for Healthcare (HEALTH), 3(4): 47:1–47:??, October 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3532628.

Adhikary:2023:SML

[ALF^{+}23] Rishiraj Adhikary, Dhruvi Lodhavia, Chris Francis, Rohit Patil, Tanmay Srivastava, Prerna Khanna, Nipun Batra, Joseph Breda, Jacob Peplinski, and Shwetak Patel. SpiroMask: Measuring lung function using consumer-grade masks. ACM Transactions on Computing for Healthcare (HEALTH), 4(1): 9:1–9:??, January 2023. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3570167.

Amon:2022:FVR

[AMN^{+}22] Mary Jean Amon, Stephen Mattingly, Aaron Necaise, Gloria Mark, Nitesh Chawla, Anind Dey, and Sidney D’mello. Flexibility versus routineness in multimodal health indicators: a sensor-based longitudinal in situ study of information workers. ACM Transactions on Computing for Healthcare (HEALTH), 3(3): 36:1–36:??, July 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3514259.

Aggarwal:2020:PDU

[APH^{+}20] Deepti Aggarwal, Bernd Ploderer, Thuong Hoang, Frank Vetere, and Mark Bradford. Physiotherapy over a distance: The use of wearable technology for video consultations in hospital settings. ACM Transactions on Computing for Healthcare (HEALTH), 1(4):21:1–21:29, December 2020. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3383305.

Akbari:2023:SBS

[ASA23] Fateme Akbari, Kamran Sartipi, and Norm Archer. Synthetic behavior sequence generation using generative adversarial networks. ACM Transactions on Computing for Healthcare (HEALTH), 4(1):7:1–7:??, January 2023. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3563950.

Ali:2021:AHP

[ASL^{+}21] Mohammad Rafayet Ali, Taylan Sen, Qianyi Li, Raina Langevin, Taylor Myers, E. Ray Dorsey, Saloni Sharma, and Ehsan Hoque. Analyzing head pose in remotely collected videos of people with Parkinson’s disease. ACM Transactions on Computing for Healthcare (HEALTH), 2(4): 30:1–30:13, October 2021. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3459669.

Bertini:2022:ASC

[BAL^{+}22] Flavio Bertini, Davide Allevi, Gianluca Lutero, Danilo Montesi, and Laura Calzà. Automatic speech classifier for mild cognitive impairment and early dementia. ACM Transactions on Computing for Healthcare (HEALTH), 3(1):8:1–8:11, January 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3469089.

Baez:2022:AEN

[BBMD^{+}22] Pablo Báez, Felipe Bravo-Marquez, Jocelyn Dunstan, Matías Rojas, and Fabián Villena. Automatic extraction of nested entities in clinical referrals in Spanish. ACM Transactions on Computing for Healthcare (HEALTH), 3(3):28:1– 28:??, July 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3498324.

Bardram:2023:DPH

[BCPM^{+}23] Jakob E. Bardram, Claus Cramer-Petersen, Alban Maxhuni, Mads V. S. Christensen, Per Bækgaard, Dan R. Persson, Nanna Lind, Merete B. Christensen, Kirsten Nørgaard, Jayden Khakurel, Timothy C. Skinner, Dagmar Kownatka, and Allan Jones. DiaFocus: a personal health technology for adaptive assessment in longterm management of Type 2 diabetes. ACM Transactions on Computing for Healthcare (HEALTH), 4(2):13:1–13:??, April 2023. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3586579.

Bellio:2021:OBA

[BFO^{+}21] Maura Bellio, Dominic Furniss, Neil P. Oxtoby, Sara Garbarino, Nicholas C. Firth, Annemie Ribbens, Daniel C. Alexander, and Ann Blandford. Opportunities and barriers for adoption of a decisionsupport tool for Alzheimer’s disease. ACM Transactions on Computing for Healthcare (HEALTH), 2(4):32:1–32:19, October 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3462764.

Bogdan:2022:ISI

[BGL22] Paul Bogdan, Radu Grosu, and Insup Lee. Introduction to the special issue on Internet-of-Medical-Things. ACM Transactions on Computing for Healthcare (HEALTH), 3(3): 26:1–26:??, July 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3547656.

Bhatia:2020:IIS

[BKS20] Munish Bhatia, Simranpreet Kaur, and Sandeep K. Sood. IoT-inspired smart toilet system for home-based urine infection prediction. ACM Transactions on Computing for Healthcare (HEALTH), 1(3): 14:1–14:25, July 2020. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3379506.

Morshed:2022:FMC

[BKS^{+}22] Mehrab Bin Morshed, Samruddhi Shreeram Kulkarni, Koustuv Saha, Richard Li, Leah G. Roper, Lama Nachman, Hong Lu, Lucia Mirabella, Sanjeev Srivastava, Kaya de Barbaro, Munmun de Choudhury, Thomas Plötz, and Gregory Abowd. Food, mood, context: Examining college students’ eating context and mental wellbeing. ACM Transactions on Computing for Healthcare (HEALTH), 3(4):38:1–38:??, October 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3533390.

BessaDaCosta:2021:PEP

[BMC^{+}21] Anderson Bessa Da Costa, Larissa Moreira, Daniel Ciampi De Andrade, Adriano Veloso, and Nivio Ziviani. Predicting the evolution of pain relief: Ensemble learning by diversifying model explanations. ACM Transactions on Computing for Healthcare (HEALTH), 2(4): 36:1–36:28, October 2021. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3466781.

Banerjee:2021:BNE

[BPDB21] Pratyay Banerjee, Kuntal Kumar Pal, Murthy Devarakonda, and Chitta Baral. Biomedical named entity recognition via knowledge guidance and question answering. ACM Transactions on Computing for Healthcare (HEALTH), 2(4): 33:1–33:24, October 2021. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3465221.

Burrello:2022:ETC

[BPR^{+}22] Alessio Burrello, Daniele Jahier Pagliari, Pierangelo Maria Rapa, Matilde Semilia, Matteo Risso, Tommaso Polonelli, Massimo Poncino, Luca Benini, and Simone Benatti. Embedding temporal convolutional networks for energy-efficient PPG-based heart rate monitoring. ACM Transactions on Computing for Healthcare (HEALTH), 3(2):19:1–19:25, April 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487910.

Bellot:2020:FML

[BV20] Alexis Bellot and Mihaela Van Der Schaar. Flexible modelling of longitudinal medical data: a Bayesian nonparametric approach. ACM Transactions on Computing for Healthcare (HEALTH), 1(1): 3:1–3:15, March 2020. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3377164.

Beh:2023:MAP

[BYL^{+}23] Win-Ken Beh, Yu-Chia Yang, Yi-Cheng Lo, Yun-Chieh Lee, and An-Yeu(Andy) Wu. Machine-aided PPG signal quality assessment (SQA) for multi-mode physiological signal monitoring. ACM Transactions on Computing for Healthcare (HEALTH), 4(2): 14:1–14:??, April 2023. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3587256.

Constantinides:2023:SUP

[CBF^{+}23] Argyris Constantinides, Marios Belk, Christos Fidas, Roy Beumers, David Vidal, Wanting Huang, Juliana Bowles, Thais Webber, Agastya Silvina, and Andreas Pitsillides. Security and usability of a personalized user authentication paradigm: Insights from a longitudinal study with three healthcare organizations. ACM Transactions on Computing for Healthcare (HEALTH), 4(1): 2:1–2:??, January 2023. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3564610.

Chen:2021:FED

[cJCCGW21] Kun chih (Jimmy) Chen, Po-Chen Chien, Zi-Jie Gao, and Chi-Hsun Wu. A fast ECG diagnosis by using non-uniform spectral analysis and the artificial neural network. ACM Transactions on Computing for Healthcare (HEALTH), 2(3): 24:1–24:21, July 2021. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3453174.

Chikka:2020:CCM

[CK20] Veera Raghavendra Chikka and Kamalakar Karlapalem. Conformance checking methodology across discharge summaries and standard treatment guidelines. ACM Transactions on Computing for Healthcare (HEALTH), 1(3):13:1– 13:19, July 2020. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3377328.

Chuan:2021:CEC

[CM21] Ching-Hua Chuan and Susan Morgan. Creating and evaluating chatbots as eligibility assistants for clinical trials: an active deep learning approach towards user-centered classification. ACM Transactions on Computing for Healthcare (HEALTH), 2(1):6:1–6:19, January 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3403575.

Chaudhuri:2024:GAC

[CMS24] Anik Chaudhuri, Anwoy Kumar Mohanty, and Manoranjan Satpathy. GPU acceleration of a conjugate exponential model for cancer tissue heterogeneity. ACM Transactions on Computing for Healthcare (HEALTH), 5(2):7:1–7:??, April 2024. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3643710.

Crovari:2022:GCA

[CPP^{+}22] Pietro Crovari, Sara Pidò, Pietro Pinoli, Anna Bernasconi, Arif Canakoglu, Franca Garzotto, and Stefano Ceri. GeCo-Agent: a conversational agent for empowering genomic data extraction and analysis. ACM Transactions on Computing for Healthcare (HEALTH), 3(1): 3:1–3:29, January 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3464383.

Cook:2022:DSB

[CSSE22] Diane J. Cook, Miranda Strickland, and Maureen Schmitter-Edgecombe. Detecting smartwatch-based behavior change in response to a multi-domain brain health intervention. ACM Transactions on Computing for Healthcare (HEALTH), 3(3):33:1– 33:??, July 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3508020.

Chen:2021:ECE

[CWR^{+}21] Dawei Chen, Jiangtao Wang, Wenjie Ruan, Qiang Ni, and Sumi Helal. Enabling costeffective population health monitoring by exploiting spatiotemporal correlation: an empirical study. ACM Transactions on Computing for Healthcare (HEALTH), 2(2): 11:1–11:19, April 2021. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3428665.

Chen:2023:QGC

[CWT23] Long Chen, Jiangtao Wang, and Piyushimita (Vonu) Thakuriah. Quality-guaranteed and costeffective population health profiling: a deep active learning approach. ACM Transactions on Computing for Healthcare (HEALTH), 4(4):22:1–22:??, October 2023. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617179.

Dedhia:2023:SSC

[DBJ23] Bhishma Dedhia, Roshini Balasubramanian, and Niraj K. Jha. SCouT: Synthetic counterfactuals via spatiotemporal transformers for actionable healthcare. ACM Transactions on Computing for Healthcare (HEALTH), 4(4):23:1–23:??, October 2023. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3617180.

Demirel:2023:DDE

[DCA23] Berken Utku Demirel, Luke Chen, and Mohammad Abdullah Al Faruque. Data-driven energy-efficient adaptive sampling using deep reinforcement learning. ACM Transactions on Computing for Healthcare (HEALTH), 4(3):19:1–19:??, July 2023. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3598301.

Ding:2022:ABU

[DL22] Haoran Ding and Xiao Luo. Attention-based unsupervised keyphrase extraction and phrase graph for COVID-19 medical literature retrieval. ACM Transactions on Computing for Healthcare (HEALTH), 3(1):12:1–12:16, January 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473939.

Darco:2024:ASI

[DMWZ24] Luigi D’arco, Graham Mccalmont, Haiying Wang, and Huiru Zheng. Application of smart insoles for recognition of activities of daily living: a systematic review. ACM Transactions on Computing for Healthcare (HEALTH), 5(1):4:1–4:??, January 2024. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3633785.

Devarakonda:2022:ISI

[DV22] Murthy V. Devarakonda and Ellen M. Voorhees. Introduction to the special issue on computational methods for biomedical NLP. ACM Transactions on Computing for Healthcare (HEALTH), 3(1): 1:1–1:2, January 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3492302.

Elor:2020:SSC

[EPM^{+}20] Aviv Elor, Michael Powell, Evanjelin Mahmoodi, Nico Hawthorne, Mircea Teodorescu, and Sri Kurniawan. On shooting stars: Comparing CAVE and HMD immersive virtual reality exergaming for adults with mixed ability. ACM Transactions on Computing for Healthcare (HEALTH), 1(4):22:1–22:22, December 2020. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3396249.

Elor:2022:BIM

[EWKM22] Aviv Elor, Steve Whittaker, Sri Kurniawan, and Sam Michael. BioLumin: an immersive mixed reality experience for interactive microscopic visualization and biomedical research annotation. ACM Transactions on Computing for Healthcare (HEALTH), 3(4): 44:1–44:??, October 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3548777.

Fagert:2022:CVH

[FBS^{+}22] Jonathon Fagert, Amelie Bonde, Sruti Srinidhi, Sarah Hamilton, Pei Zhang, and Hae Young Noh. Clean vibes: Hand washing monitoring using structural vibration sensing. ACM Transactions on Computing for Healthcare (HEALTH), 3(3):34:1– 34:??, July 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3511890.

Fu:2022:SLR

[FHS22] Yu Fu, Yan Hu, and Veronica Sundstedt. A systematic literature review of virtual, augmented, and mixed reality game applications in healthcare. ACM Transactions on Computing for Healthcare (HEALTH), 3(2):22:1–22:27, April 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3472303.

Fan:2024:CPF

[FJK24] Shaojing Fan, Ramesh C. Jain, and Mohan S. Kankanhalli. A comprehensive picture of factors affecting user willingness to use mobile health applications. ACM Transactions on Computing for Healthcare (HEALTH), 5(1):2:1–2:??, January 2024. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3626962.

Feiger:2022:EUN

[FLSC^{+}22] Bradley Feiger, Erick Lorenzana-Saldivar, Colin Cooke, Roarke Horstmeyer, Muath Bishawi, Julie Doberne, G. Chad Hughes, David Ranney, Soraya Voigt, and Amanda Randles. Evaluation of U-net based architectures for automatic aortic dissection segmentation. ACM Transactions on Computing for Healthcare (HEALTH), 3(1):11:1–11:16, January 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3472302.

Feng:2023:TSC

[FWWC23] Yujie Feng, Jiangtao Wang, Yasha Wang, and Xu Chu. Towards sustainable compressive population health: a GANbased year-by-year imputation method. ACM Transactions on Computing for Healthcare (HEALTH), 4(1):8:1–8:??, January 2023. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571159.

Ghosal:2022:IHA

[GDUW22] Sagnik Ghosal, Debanjan Das, Venkanna Udutalapally, and Preetam Narayan Wasnik. iNAP: a hybrid approach for NonInvasive anemia-polycythemia detection in the IoMT. ACM Transactions on Computing for Healthcare (HEALTH), 3(3): 31:1–31:??, July 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3503466.

Goldstein:2020:TFF

[GKKS20] Orpaz Goldstein, Mohammad Kachuee, Kimmo Karkkainen, and Majid Sarrafzadeh. Targetfocused feature selection using uncertainty measurements in healthcare data. ACM Transactions on Computing for Healthcare (HEALTH), 1(3): 15:1–15:17, July 2020. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3383685.

Grohn:2023:QMB

[GLH^{+}23] Tommi Gröhn, Sammeli Liikkanen, Teppo Huttunen, Mika Mäkinen, Pasi Liljeberg, and Pekka Marttinen. Quantifying movement behavior of chronic low back pain patients in virtual reality. ACM Transactions on Computing for Healthcare (HEALTH), 4(2): 11:1–11:??, April 2023. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3582487.

Grani:2022:SPH

[GLV22] Giorgio Grani, Andrea Lenzi, and Paola Velardi. Supporting personalized health care with social media analytics: an application to hypothyroidism. ACM Transactions on Computing for Healthcare (HEALTH), 3(1):4:1–4:28, January 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3468781.

Goh:2024:WTW

[GP24] Teck Lun Goh and Li-Shiuan Peh. WalkingWizard — a truly wearable EEG headset for everyday use. ACM Transactions on Computing for Healthcare (HEALTH), 5(2):9:1–9:??, April 2024. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3648106.

Gupta:2022:OPE

[GPBB22] Mehak Gupta, Thao-Ly T. Phan, H. Timothy Bunnell, and Rahmatollah Beheshti. Obesity prediction with EHR data: a deep learning approach with interpretable elements. ACM Transactions on Computing for Healthcare (HEALTH), 3(3):32:1–32:??, July 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3506719.

Gaddipati:2022:GAF

[GS22] Divya Jyothi Gaddipati and Jayanthi Sivaswamy. Glaucoma assessment from fundus images with fundus to OCT feature space mapping. ACM Transactions on Computing for Healthcare (HEALTH), 3(1): 10:1–10:15, January 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3470979.

Gao:2022:ERR

[GSG^{+}22] Ye Gao, Asif Salekin, Kristina Gordon, Karen Rose, Hongning Wang, and John Stankovic. Emotion recognition robust to indoor environmental distortions and non-targeted emotions using out-of-distribution detection. ACM Transactions on Computing for Healthcare (HEALTH), 3(2):15:1–15:22, April 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3492300.

Gu:2022:DSL

[GTC^{+}22] Yu Gu, Robert Tinn, Hao Cheng, Michael Lucas, Naoto Usuyama, Xiaodong Liu, Tristan Naumann, Jianfeng Gao, and Hoifung Poon. Domainspecific language model pretraining for biomedical natural language processing. ACM Transactions on Computing for Healthcare (HEALTH), 3(1): 2:1–2:23, January 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3458754.

Harris:2021:FGS

[HCZ21] Jonathan J. Harris, Ching-Hua Chen, and Mohammed J. Zaki. A framework for generating summaries from temporal personal health data. ACM Transactions on Computing for Healthcare (HEALTH), 2(3): 21:1–21:43, July 2021. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3448672.

Hurley:2023:CPO

[HDD^{+}23] Nathan C. Hurley, Sanket S.

Dhruva, Nihar R. Desai, Joseph R. Ross, Che G. Ngufor, Frederick Masoudi, Harlan M. Krumholz, and Bobak J. Mortazavi. Clinical phenotyping with an outcomes-driven mixture of experts for patient matching and risk estimation. ACM Transactions on Computing for Healthcare (HEALTH), 4(4):21:1– 21:??, October 2023. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3616021.

Hussein:2020:ADA

[HDM^{+}20] Amir Hussein, Marc Djandji, Reem A. Mahmoud, Mohamad Dhaybi, and Hazem Hajj. Augmenting DL with adversarial training for robust prediction of epilepsy seizures. ACM Transactions on Computing for Healthcare (HEALTH), 1(3): 18:1–18:18, July 2020. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3386580.

Hu:2023:ICB

[HK23] Hengyi Hu and Larry Kerschberg. Improving causal Bayesian networks using expertise in authoritative medical ontologies. ACM Transactions on Computing for Healthcare (HEALTH), 4(4):20:1–20:??, October 2023. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3604561.

Herrlich:2020:PNT

[HRW^{+}20] Marc Herrlich, Anke V. Reinschluessel, Markus Willems, Nils Langhorst, David Black, Tanja Döring, Christian Rieder, Ron Kikinis, and Rainer Malaka. Put that needle there: Customized flexible onbody thin-film displays for medical navigation. ACM Transactions on Computing for Healthcare (HEALTH), 1(3): 16:1–16:17, July 2020. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3386307.

Hurley:2021:SCO

[HSK^{+}21] Nathan C. Hurley, Erica S.

Spatz, Harlan M. Krumholz, Roozbeh Jafari, and Bobak J. Mortazavi. A survey of challenges and opportunities in sensing and analytics for risk factors of cardiovascular disorders. ACM Transactions on Computing for Healthcare (HEALTH), 2(1):9:1–9:42, January 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3417958.

Hussain:2022:NIT

[HSZ^{+}22] Zawar Hussain, Quan Z.

Sheng, Wei Emma Zhang, Jorge Ortiz, and Seyedamin Pouriyeh. Non-invasive techniques for monitoring different aspects of sleep: a comprehensive review. ACM Transactions on Computing for Healthcare (HEALTH), 3(2):24:1–24:26, April 2022. CODEN ????

ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3491245.

Jourdan:2021:PPI

[JBBF21] Theo Jourdan, Antoine Boutet, Amine Bahi, and Carole Frindel. Privacy-preserving IoT framework for activity recognition in personal healthcare monitoring. ACM Transactions on Computing for Healthcare (HEALTH), 2(1): 3:1–3:22, January 2021. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3416947.

Jin:2020:CIC

[JCG^{+}20] Zhuochen Jin, Shuyuan Cui, Shunan Guo, David Gotz, Jimeng Sun, and Nan Cao. CarePre: an intelligent clinical decision assistance system. ACM Transactions on Computing for Healthcare (HEALTH), 1(1):6:1–6:20, March 2020. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3344258.

Jain:2022:EEH

[JJU^{+}22] Seemandhar Jain, Prarthi Jain, Prabhat K. Upadhyay, Jules M. Moualeu, and Abhishek Srivastava. An energy efficient health monitoring approach with wireless body area networks. ACM Transactions on Computing for Healthcare (HEALTH), 3(3):29:1– 29:??, July 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3501773.

Joshi:2022:FLH

[JPS22] Madhura Joshi, Ankit Pal, and Malaikannan Sankarasubbu. Federated learning for healthcare domain — pipeline, applications and challenges. ACM Transactions on Computing for Healthcare (HEALTH), 3(4): 40:1–40:??, October 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3533708.

Katsimpras:2022:IEP

[KAG^{+}22] Georgios Katsimpras, Fotis Aisopos, Peter Garrard, Maria-Esther Vidal, and Georgios Paliouras. Improving early prognosis of dementia using machine learning methods. ACM Transactions on Computing for Healthcare (HEALTH), 3(3):30:1–30:??, July 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3502433.

Kyono:2023:STE

[KBQvdS23] Trent Kyono, Ioana Bica, Zhaozhi Qian, and Mihaela van der Schaar. Selecting treatment effects models for domain adaptation using causal knowledge. ACM Transactions on Computing for Healthcare (HEALTH), 4(2):15:1–15:??, April 2023. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3587695.

Kucukozer-Cavdar:2021:DRM

[KCTTM^{+}21] Seyma Kucukozer-Cavdar, Tugba Taskaya-Temizel, Abhinav Mehrotra, Mirco Musolesi, and Peter Tino. Designing robust models for behaviour prediction using sparse data from mobile sensing: a case study of office workers’ availability for well-being interventions. ACM Transactions on Computing for Healthcare (HEALTH), 2(4): 29:1–29:33, October 2021. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3458753.

Kumar:2022:LMS

[KDD^{+}22] Shashwat Kumar, Debajyoti Datta, Guimin Dong, Lihua Cai, Mehdi Boukhechba, and Laura Barnes. Leveraging mobile sensing and Bayesian change point analysis to monitor community-scale behavioral interventions: a case study on COVID-19. ACM Transactions on Computing for Healthcare (HEALTH), 3(4): 37:1–37:??, October 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3524886.

Kyono:2021:TMI

[KGV21] Trent Kyono, Fiona J. Gilbert, and Mihaela Van Der Schaar. Triage of 2D mammographic images using multi-view multitask convolutional neural networks. ACM Transactions on Computing for Healthcare (HEALTH), 2(3):26:1–26:24, July 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3453166.

Kassaie:2022:CAC

[KIT22] Besat Kassaie, Elizabeth L.

Irving, and Frank Wm. Tompa. Computer-assisted cohort identification in practice. ACM Transactions on Computing for Healthcare (HEALTH), 3(2):17:1–17:28, April 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3483411.

Kumar:2021:MWS

[KJBD21] Devender Kumar, Steven Jeuris, Jakob E. Bardram, and Nicola Dragoni. Mobile and wearable sensing frameworks for mHealth studies and applications: a systematic review. ACM Transactions on Computing for Healthcare (HEALTH), 2(1):8:1–8:28, January 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3422158.

Khan:2021:GES

[KMK^{+}21] Aftab Khan, Sebastian Mellor, Rachel King, Balazs Janko, William Harwin, R. Simon Sherratt, Ian Craddock, and Thomas Plötz. Generalized and efficient skill assessment from IMU data with applications in gymnastics and medical training. ACM Transactions on Computing for Healthcare (HEALTH), 2(1):2:1–2:21, January 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3422168.

Kumar:2022:MCA

[KMM^{+}22] Devender Kumar, Raju Maharjan, Alban Maxhuni, Helena Dominguez, Anne Frølich, and Jakob E. Bardram. mCardia: a context-aware ECG collection system for ambulatory arrhythmia screening. ACM Transactions on Computing for Healthcare (HEALTH), 3(2): 20:1–20:28, April 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3494581.

Kristiansen:2021:MLS

[KNP^{+}21] Stein Kristiansen, Konstantinos Nikolaidis, Thomas Plagemann, Vera Goebel, Gunn Marit Traaen, Britt Øverland, Lars Aakerøy, Tove-Elizabeth Hunt, Jan Pål Loennechen, Sigurd Loe Steinshamn, Christina Holt Bendz, Ole-Gunnar Anfinsen, Lars Gullestad, and Harriet Akre. Machine learning for sleep apnea detection with unattended sleep monitoring at home. ACM Transactions on Computing for Healthcare (HEALTH), 2(2):14:1–14:25, March 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3433987.

Kumar:2023:FIB

[KP23] Vijay Kumar and Kolin Paul.

Fundus imaging-based healthcare: Present and future. ACM Transactions on Computing for Healthcare (HEALTH), 4(3): 16:1–16:??, July 2023. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3586580.

Kumar:2023:DLA

[KPPA23] Vijay Kumar, Het Patel, Kolin Paul, and Shorya Azad. Deep learning-assisted retinopathy of prematurity (ROP) screening. ACM Transactions on Computing for Healthcare (HEALTH), 4(3):18:1–18:??, July 2023. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3596223.

Kotz:2020:ISI

[KX20] David Kotz and Guoliang Xing. Introduction to the special issue on the wearable technologies for smart health. ACM Transactions on Computing for Healthcare (HEALTH), 1 (4):19:1–19:2, December 2020. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3423967.

Kotz:2021:ISI

[KX21] David Kotz and Guoliang Xing. Introduction to the special issue on the wearable technologies for smart health, Part 2. ACM Transactions on Computing for Healthcare (HEALTH), 2(1):1:1–1:2, January 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3442350.

Khan:2021:PCA

[KZK^{+}21] Aftab Khan, Alexandros Zenonos, Georgios Kalogridis, Yaowei Wang, Stefanos Vatsikas, and Mahesh Sooriyabandara. Perception clusters: Automated mood recognition using a novel cluster-driven modelling system. ACM Transactions on Computing for Healthcare (HEALTH), 2(1):4:1–4:16, January 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3422819.

Luo:2022:PPE

[LBHS22] Jiaqing Luo, Ruiyu Bai, Suining He, and Kang G. Shin. Pervasive pose estimation for fall detection. ACM Transactions on Computing for Healthcare (HEALTH), 3(3):27:1– 27:??, July 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3478027.

Li:2020:MER

[LCLS20] Chun-Tung Li, Jiannong Cao, Xue Liu, and Milos Stojmenovic. mSIMPAD: Efficient and robust mining of successive similar patterns of multiple lengths in time series. ACM Transactions on Computing for Healthcare (HEALTH), 1(4): 23:1–23:19, December 2020. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3396250.

Li:2024:ERL

[LHFC^{+}24] Pei-Xuan Li, Hsun-Ping Hsieh, Yang Fan-Chiang, Ding-You Wu, and Ching-Chung Ko. Enhancing robust liver cancer diagnosis: a contrastive multi-modality learner with lightweight fusion and effective data augmentation. ACM Transactions on Computing for Healthcare (HEALTH), 5(2): 6:1–6:??, April 2024. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3639414.

Lee:2022:CEI

[LKHU22] Kahyun Lee, Mehmet Kayaalp, Sam Henry, and Özlem Uzuner. A context-enhanced de-identification system. ACM Transactions on Computing for Healthcare (HEALTH), 3(1):6:1–6:14, January 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3470980.

Li:2020:FSM

[LVW^{+}20] Dingwen Li, Jay Vaidya, Michael Wang, Ben Bush, Chenyang Lu, Marin Kollef, and Thomas Bailey. Feasibility study of monitoring deterioration of outpatients using multimodal data collected by wearables. ACM Transactions on Computing for Healthcare (HEALTH), 1(1): 5:1–5:22, March 2020. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3344256.

Le:2023:MAM

[LYH^{+}23] Hieu Hanh Le, Tatsuhiro Yamada, Yuichi Honda, Takatoshi Sakamoto, Ryosuke Matsuo, Tomoyoshi Yamazaki, Kenji Araki, and Haruo Yokota. Methods for analyzing medical-order sequence variants in sequential pattern mining for electronic medical record systems. ACM Transactions on Computing for Healthcare (HEALTH), 4(1):3:1–3:??, January 2023. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3561825.

Minot:2022:IBM

[MCM^{+}22] Joshua R. Minot, Nicholas Cheney, Marc Maier, Danne C. Elbers, Christopher M. Danforth, and Peter Sheridan Dodds. Interpretable bias mitigation for textual data: Reducing genderization in patient notes while maintaining classification performance. ACM Transactions on Computing for Healthcare (HEALTH), 3(4): 39:1–39:??, October 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3524887.

Makitalo:2020:HDM

[MFMF^{+}20] Niko Mäkitalo, Daniel Flores-Martin, Huber Flores, Eemil Lagerspetz, Francois Christophe, Petri Ihantola, Masiar Babazadeh, Pan Hui, Juan Manuel Murillo, Sasu Tarkoma, and Tommi Mikkonen. Human data model: Improving programmability of health and well-being data for enhanced perception and interaction. ACM Transactions on Computing for Healthcare (HEALTH), 1(4):26:1–26:39, December 2020. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3402524.

Marshall:2022:SHF

[MJPMC22] Jermaine Marshall, Priscilla Jimenez-Pazmino, Ronald Metoyer, and Nitesh V. Chawla. A survey on healthy food decision influences through technological innovations. ACM Transactions on Computing for Healthcare (HEALTH), 3(2): 25:1–25:27, April 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3494580.

Meyer:2020:LDC

[MKE^{+}20] Jochen Meyer, Judy Kay, Daniel A. Epstein, Parisa Eslambolchilar, and Lie Ming Tang. A life of data: Characteristics and challenges of very long term self-tracking for health and wellness. ACM Transactions on Computing for Healthcare (HEALTH), 1(2): 11:1–11:4, March 2020. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3373719.

Mauldin:2021:EDL

[MNMC21] Taylor Mauldin, Anne H. Ngu, Vangelis Metsis, and Marc E. Canby. Ensemble deep learning on wearables using small datasets. ACM Transactions on Computing for Healthcare (HEALTH), 2(1):5:1–5:30, January 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3428666.

Mahajan:2020:THW

[MPK20] Aman Mahajan, Gregory Pottie, and William Kaiser. Transformation in healthcare by wearable devices for diagnostics and guidance of treatment. ACM Transactions on Computing for Healthcare (HEALTH), 1(1):2:1–2:12, March 2020. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3361561.

Mishra:2020:CDP

[MPL^{+}20] Varun Mishra, Gunnar Pope, Sarah Lord, Stephanie Lewia, Byron Lowens, Kelly Caine, Sougata Sen, Ryan Halter, and David Kotz. Continuous detection of physiological stress with commodity hardware. ACM Transactions on Computing for Healthcare (HEALTH), 1(2): 8:1–8:30, April 2020. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3361562.

Malavazos:2021:NRH

[MPMP21] Konstantinos Malavazos, Maria Papadogiorgaki, Pavlos Malakonakis, and Ioannis Papaefstathiou. Novel reconfigurable hardware systems for tumor growth prediction. ACM Transactions on Computing for Healthcare (HEALTH), 2(4):

28:1–28:27, October 2021. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3454126.

Nikolaidis:2022:MHS

[NKP^{+}22] Konstantinos Nikolaidis, Stein Kristiansen, Thomas Plagemann, Vera Goebel, Knut Liestøl, Mohan Kankanhalli, Gunn-Marit Traaen, Britt Øverland, Harriet Akre, Lars Aakerøy, and Sigurd Steinshamn. My Health Sensor, My Classifier — adapting a trained classifier to unlabeled enduser data. ACM Transactions on Computing for Healthcare (HEALTH), 3(4):48:1–48:??, October 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3559767.

Nosakhare:2020:TAR

[NP20] Ehimwenma Nosakhare and Rosalind Picard. Toward assessing and recommending combinations of behaviors for improving health and wellbeing. ACM Transactions on Computing for Healthcare (HEALTH), 1(1):4:1–4:29, March 2020. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3368958.

Naeini:2023:DLB

[NSA^{+}23] Emad Kasaeyan Naeini, Fatemeh Sarhaddi, Iman Azimi, Pasi Liljeberg, Nikil Dutt, and Amir M. Rahmani. A deep learning-based PPG quality assessment approach for heart rate and heart rate variability. ACM Transactions on Computing for Healthcare (HEALTH), 4(4):24:1–24:??, October 2023. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3616019.

Newaz:2021:SSP

[NSRU21] Akm Iqtidar Newaz, Amit Kumar Sikder, Mohammad Ashiqur Rahman, and A. Selcuk Uluagac. A survey on security and privacy issues in modern healthcare systems: Attacks and defenses. ACM Transactions on Computing for Healthcare (HEALTH), 2(3):27:1– 27:44, July 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3453176.

Pourebadi:2022:FEM

[PR22] Maryam Pourebadi and Laurel D. Riek. Facial expression modeling and synthesis for patient simulator systems: Past, present, and future. ACM Transactions on Computing for Healthcare (HEALTH), 3(2): 23:1–23:32, April 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3483598.

Pugh:2022:EAC

[PRB^{+}22] Sydney Pugh, Ivan Ruchkin, Christopher Bonafide, Sara Demauro, Oleg Sokolsky, Insup Lee, and James Weimer. Evaluating alarm classifiers with high-confidence data programming. ACM Transactions on Computing for Healthcare (HEALTH), 3(4):43:1–43:??, October 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3549942.

Reichert:2021:SAC

[RBS21] Leonie Reichert, Samuel Brack, and BjÖRN Scheuermann. A survey of automatic contact tracing approaches using Bluetooth low energy. ACM Transactions on Computing for Healthcare (HEALTH), 2(2): 18:1–18:33, March 2021. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3444847.

Rohani:2021:BUA

[RFJKB21] Darius A. Rohani, Maria Faurholt-Jepsen, Lars V. Kessing, and Jakob E. Bardram. Benefits of using activity recommender technology for self-management of depressive symptoms. ACM Transactions on Computing for Healthcare (HEALTH), 2(4):31:1–31:21, October 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3462212.

Rostaminia:2022:PRS

[RHK^{+}22] Soha Rostaminia, Seyedeh Zohreh Homayounfar, Ali Kiaghadi, Trisha Andrew, and Deepak Ganesan. PhyMask: Robust sensing of brain activity and physiological signals during sleep with an all-textile eye mask. ACM Transactions on Computing for Healthcare (HEALTH), 3(3):35:1– 35:??, July 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3513023.

Rissola:2021:SCM

[RLC21] Esteban A. Ríssola, David E.

Losada, and Fabio Crestani. A survey of computational methods for online mental state assessment on social media. ACM Transactions on Computing for Healthcare (HEALTH), 2(2):17:1–17:31, March 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3437259.

Rahman:2022:MMT

[RM22] Md Juber Rahman and Bashir I.

Morshed. A minimalist method toward severity assessment and progression monitoring of obstructive sleep apnea on the edge. ACM Transactions on Computing for Healthcare (HEALTH), 3(2):16:1–16:16, April 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3479432.

Santini:2021:OCV

[San21]  Simone Santini.  Optimal

COVID-19 vaccination strategies with limited vaccine and delivery capabilities: a model based on social networks. ACM Transactions on Computing for Healthcare (HEALTH), 2(4): 34:1–34:16, October 2021. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3466622.

Sun:2022:CDE

[SAW22] Ting-Wei Sun, Danish Ali, and Ayeu (Andy) Wu. Compresseddomain ECG-based biometric user identification using task-driven dictionary learning. ACM Transactions on Computing for Healthcare (HEALTH), 3(3):26:1–26:??, July 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3461701.

Saffarpour:2023:PPA

[SBR^{+}23] Mahya Saffarpour, Debraj Basu, Fatemeh Radaei, Kourosh Vali, Jason Y. Adams, Chen-Nee Chuah, and Soheil Ghiasi. Physiowise: a physicsaware approach to dicrotic notch identification. ACM Transactions on Computing for Healthcare (HEALTH), 4(2): 10:1–10:??, April 2023. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3578556.

Shojaee:2022:AWT

[SCJ22] Parshin Shojaee, Xiaoyu Chen, and Ran Jin. Adaptively weighted top-N recommendation for organ matching. ACM Transactions on Computing for Healthcare (HEALTH), 3(1): 9:1–9:29, January 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3469657.

Shi:2023:IAP

[SCL^{+}23] Keqin Shi, Zhen Chen, Xuejing Li, Zhifei Xiao, Weiqiang Sun, and Weisheng Hu. Inferring activity patterns from sparse step counts data with recurrent neural networks. ACM Transactions on Computing for Healthcare (HEALTH), 4(1): 4:1–4:??, January 2023. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3560468.

Sprint:2022:MFS

[SCSEH22] Gina Sprint, Diane J. Cook, Maureen Schmitter-Edgecombe, and Lawrence B. Holder. Multimodal fusion of smart home and text-based behavior markers for clinical assessment prediction. ACM Transactions on Computing for Healthcare (HEALTH), 3(4):41:1–41:??, October 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3531231.

Sun:2021:ABD

[SDS^{+}21] Zhaohong Sun, Wei Dong, Jinlong Shi, Kunlun He, and Zhengxing Huang. Attentionbased deep recurrent model for survival prediction. ACM Transactions on Computing for Healthcare (HEALTH), 2(4): 35:1–35:18, October 2021. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3466782.

Slijepcevic:2022:EML

[SHL^{+}22] Djordje Slijepcevic, Fabian Horst, Sebastian Lapuschkin, Brian Horsak, Anna-Maria Raberger, Andreas Kranzl, Wojciech Samek, Christian Breiteneder, Wolfgang Immanuel Schöllhorn, and Matthias Zeppelzauer. Explaining machine learning models for clinical gait analysis. ACM Transactions on Computing for Healthcare (HEALTH), 3(2): 14:1–14:27, April 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3474121.

Sharma:2020:IWR

[SJMH20] Surya Sharma, Phillip Jasper, Eric Muth, and Adam Hoover. The impact of walking and resting on wrist motion for automated detection of meals. ACM Transactions on Computing for Healthcare (HEALTH), 1(4):24:1–24:19, December 2020. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3407623.

Stankovic:2020:IIE

[SL20] John A. Stankovic and Insup Lee. Inaugural issue editorial. ACM Transactions on Computing for Healthcare (HEALTH), 1(1):1:1, March 2020. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3375711.

Song:2024:LRS

[SLT^{+}24] Siyang Song, Yiming Luo, Tugba Tumer, Changzeng Fu, Michel Valstar, and Hatice Gunes. Loss relaxation strategy for noisy facial video-based automatic depression recognition. ACM Transactions on Computing for Healthcare (HEALTH), 5(2):12:1–12:??, April 2024. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3648696.

Sun:2020:HPT

[SMSV20] Xuetong Sun, Sarah B. Murthi, Gary Schwartzbauer, and Amitabh Varshney. Highprecision 5DoF tracking and visualization of catheter placement in EVD of the brain using AR. ACM Transactions on Computing for Healthcare (HEALTH), 1(2):9:1–9:18, March 2020. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3365678.

Shojaati:2022:DCM

[SO22] Narjes Shojaati and Nathaniel D. Osgood. Dynamic computational models and simulations of the opioid crisis: a comprehensive survey. ACM Transactions on Computing for Healthcare (HEALTH), 3(1): 13:1–13:25, January 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3469658.

Senaratne:2022:CRM

[SOEM22] Hashini Senaratne, Sharon Oviatt, Kirsten Ellis, and Glenn Melvin. A critical review of multimodal-multisensor analytics for anxiety assessment. ACM Transactions on Computing for Healthcare (HEALTH), 3(4):46:1–46:??, October 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3556980.

Sweeney:2021:CCH

[SPE^{+}21] Colm Sweeney, Courtney Potts, Edel Ennis, Raymond Bond, Maurice D. Mulvenna, Siobhan O’neill, Martin Malcolm, Lauri Kuosmanen, Catrine Kostenius, Alex Vakaloudis, Gavin Mcconvey, Robin Turkington, David Hanna, Heidi Nieminen, Anna-Kaisa Vartiainen, Alison Robertson, and Michael F. Mctear. Can chatbots help support a Person’s mental health? perceptions and views from mental healthcare professionals and experts. ACM Transactions on Computing for Healthcare (HEALTH), 2(3): 25:1–25:15, July 2021. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3453175.

Shen:2022:MAH

[SQHS22] Cong Shen, Zhaozhi Qian, Alihan Huyuk, and Mihaela Van Der Schaar. MARS: Assisting human with information processing tasks using machine learning. ACM Transactions on Computing for Healthcare (HEALTH), 3(2): 21:1–21:19, April 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3494582.

Thimbleby:2022:NNO

[Thi22]  Harold Thimbleby.  NHS

number open source software: Implications for digital health regulation and development. ACM Transactions on Computing for Healthcare (HEALTH), 3(4):42:1–42:??, October 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3538382.

Thambawita:2020:ESC

[TJH^{+}20] Vajira Thambawita, Debesh Jha, Hugo Lewi Hammer, Håvard D. Johansen, Dag Johansen, Pål Halvorsen, and Michael A. Riegler. An extensive study on cross-dataset bias and evaluation metrics interpretation for machine learning applied to gastrointestinal tract abnormality classification. ACM Transactions on Computing for Healthcare (HEALTH), 1(3):17:1–17:29, July 2020. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3386295.

Tran:2021:AGG

[TKK21] Tung Tran, Ramakanth Kavuluru, and Halil Kilicoglu. Attention-gated graph convolutions for extracting drug interaction information from drug labels. ACM Transactions on Computing for Healthcare (HEALTH), 2(2):10:1–10:19, March 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3423209.

Tazi:2024:SAP

[TND^{+}24] Faiza Tazi, Archana Nandakumar, Josiah Dykstra, Prashanth Rajivan, and Sanchari Das. SoK: Analyzing privacy and security of healthcare data from the user perspective. ACM Transactions on Computing for Healthcare (HEALTH), 5(2):11:1–11:??, April 2024. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3650116.

Thakur:2024:SRH

[TP24] Dipanwita Thakur and Arindam Pal. Subsampled randomized Hadamard transformationbased ensemble extreme learning machine for human activity recognition. ACM Transactions on Computing for Healthcare (HEALTH), 5(1):5:1–5:??, January 2024. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3634813.

Tariq:2022:BGB

[TVDB22] Amara Tariq, Marly Van Assen, Carlo N. De Cecco, and Imon Banerjee. Bridging the gap between structured and free-form radiology reporting: a case-study on coronary CT angiography. ACM Transactions on Computing for Healthcare (HEALTH), 3(1):7:1–7:20, January 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3474831.

vanBerkel:2021:DVM

[vBAS^{+}21] Niels van Berkel, Omer F. Ahmad, Danail Stoyanov, Laurence Lovat, and Ann Blandford. Designing visual markers for continuous artificial intelligence support: a colonoscopy case study. ACM Transactions on Computing for Healthcare (HEALTH), 2(1):7:1–7:24, January 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3422156.

Vargemidis:2020:WPA

[VGS^{+}20] Dimitri Vargemidis, Kathrin Gerling, Katta Spiel, Vero Vanden Abeele, and Luc Geurts. Wearable physical activity tracking systems for older adults — a systematic review. ACM Transactions on Computing for Healthcare (HEALTH), 1(4):25:1–25:37, December 2020. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3402523.

Verreydt:2021:SPR

[VYJ21] Stef Verreydt, Koen Yskout, and Wouter Joosen. Security and privacy requirements for electronic consent: a systematic literature review. ACM Transactions on Computing for Healthcare (HEALTH), 2(2): 16:1–16:24, March 2021. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3433995.

Wang:2021:SSB

[WHYM21] Xuyu Wang, Runze Huang, Chao Yang, and Shiwen Mao. Smartphone sonar-based contact-free respiration rate monitoring. ACM Transactions on Computing for Healthcare (HEALTH), 2(2):15:1– 15:26, March 2021. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3436822.

Wu:2021:EDH

[WKS^{+}21] Wenlong Wu, James M.

Keller, Marjorie Skubic, Mihail Popescu, and Kari R. Lane. Early detection of health changes in the elderly using in-home multi-sensor data streams. ACM Transactions on Computing for Healthcare (HEALTH), 2(3):22:1–22:23, July 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3448671.

Woodward:2024:CDL

[WKT24] Kieran Woodward, Eiman Kanjo, and Athanasios Tsanas. Combining deep learning with signal-image encoding for multi-modal mental wellbeing classification. ACM Transactions on Computing for Healthcare (HEALTH), 5(1):3:1–3:??, January 2024. CODEN ????

ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3631618.

Wan:2021:SBE

[WMKS21] Cheng Wan, Andrew W. Mchill, Elizabeth B. Klerman, and Akane Sano. Sensor-based estimation of dim light melatonin onset using features of two time scales. ACM Transactions on Computing for Healthcare (HEALTH), 2(3):20:1– 20:15, July 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3447516.

Wang:2021:CPP

[WOM^{+}21] Chongyang Wang, Temitayo A. Olugbade, Akhil Mathur, Amanda C. DE C. Williams, Nicholas D. Lane, and Nadia Bianchi-Berthouze. Chronic pain protective behavior detection with deep learning. ACM Transactions on Computing for Healthcare (HEALTH), 2(3): 23:1–23:24, July 2021. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3449068.

Wang:2023:AAB

[WWF^{+}23] Yiyuan Wang, Guillaume Wattelez, Stéphane Frayon, Corinne Caillaud, Olivier Galy, and Kalina Yacef. ABIPA: ARIMA-based integration of accelerometer-based physical activity for adolescent weight status prediction. ACM Transactions on Computing for Healthcare (HEALTH), 4(1): 6:1–6:??, January 2023. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3561611.

Wang:2020:CBV

[WYM20] Xuyu Wang, Chao Yang, and Shiwen Mao. On CSI-based vital sign monitoring using commodity WiFi. ACM Transactions on Computing for Healthcare (HEALTH), 1(3):12:1– 12:27, July 2020. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3377165.

Wang:2020:EAE

[WYZ^{+}20] Zhu Wang, Zhiwen Yu, Bobo Zhao, Bin Guo, Chao Chen, and Zhiyong Yu. Emotion-Sense: an adaptive emotion recognition system based on wearable smart devices. ACM Transactions on Computing for Healthcare (HEALTH), 1(4): 20:1–20:17, December 2020. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3384394.

Xie:2022:PCA

[XZC^{+}22] Zongxing Xie, Bing Zhou, Xi Cheng, Elinor Schoenfeld, and Fan Ye. Passive and context-aware in-home vital signs monitoring using colocated UWB-depth sensor fusion. ACM Transactions on Computing for Healthcare (HEALTH), 3(4):45:1–45:??, October 2022. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3549941.

Yao:2024:SSR

[YCR^{+}24] Hao-Ren Yao, Nairen Cao, Katina Russell, Der-Chen Chang, Ophir Frieder, and Jeremy T. Fineman. Selfsupervised representation learning on electronic health records with graph kernel infomax. ACM Transactions on Computing for Healthcare (HEALTH), 5(2):10:1–10:??, April 2024. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3648695.

Yao:2024:ITA

[YCW^{+}24] Zhenjie Yao, Yixin Chen, Jinwei Wang, Junjuan Li, Shuohua Chen, Shouling Wu, Yanhui Tu, Ming-Hui Zhao, and Luxia Zhang. Interpretable trend analysis neural networks for longitudinal data analysis. ACM Transactions on Computing for Healthcare (HEALTH), 5(2):8:1–8:??, April 2024. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3648105.

Yang:2023:TGT

[YLZ23] Zhenyu Yang, Yantao Li, and Gang Zhou. TS-GAN: Time-series GAN for sensorbased health data augmentation. ACM Transactions on Computing for Healthcare (HEALTH), 4(2):12:1–12:??, April 2023. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3583593.

Yarlagadda:2021:SAC

[YMG^{+}21] Sri Kalyan Yarlagadda, Daniel Mas Montserrat, David Güera, Carol J. Boushey, Deborah A. Kerr, and Fengqing Zhu. Saliency-aware classagnostic food image segmentation. ACM Transactions on Computing for Healthcare (HEALTH), 2(3):19:1–19:17, July 2021. CODEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3440274.

Yfantidou:2023:YST

[YSV23] Sofia Yfantidou, Pavlos Sermpezis, and Athena Vakali. 14 years of self-tracking technology for mHealth-literature review: Lessons learned and the PAST SELF framework. ACM Transactions on Computing for Healthcare (HEALTH), 4(3): 17:1–17:??, July 2023. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

URL https://dl.acm.org/
doi/10.1145/3592621.

Yom-Tov:2020:SCU

[YT20] Elad Yom-Tov. Screening for cancer using a learning Internet advertising system. ACM Transactions on Computing for Healthcare (HEALTH), 1(2): 10:1–10:13, March 2020. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3373720.

Zargari:2023:ANA

[ZAK^{+}23] Amir Hosein Afandizadeh Zargari, Seyed Amir Hossein Aqajari, Hadi Khodabandeh, Amir Rahmani, and Fadi Kurdahi. An accurate non-accelerometer-based PPG motion artifact removal technique using CycleGAN. ACM Transactions on Computing for Healthcare (HEALTH), 4(1): 1:1–1:??, January 2023. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3563949.

Zhang:2022:APG

[ZSL^{+}22] Junqian Zhang, Yingming Sun, Hongen Liao, Jian Zhu, and Yuan Zhang. Automatic parotid gland segmentation in MVCT using deep convolutional neural networks. ACM Transactions on Computing for Healthcare (HEALTH), 3(2): 18:1–18:15, April 2022. CO-DEN ???? ISSN 2691-1957 (print), 2637-8051 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3485278.